Quota. Determine the SWL.
D: P = 840 - 3Q
S: P = 240 + 3Q
Quota: Q = 80.
1
Expert's answer
2016-05-26T12:42:03-0400
D: P = 840 - 3Q
S: P = 240 + 3Q
In equilibrium quantity demanded equals quantity supplied at equilibrium price, so:
840 - 3Q = 240 + 3Q,
6Q = 600,
Q = 100 units.
P = 840 - 3*100 = $540.
If there is a quota of Q = 80, then the price of demand is: P = 840 - 3*80 = $600,
the price of supply is: P = 240 + 3*80 = $480.
So, the deadweight loss is: DWL = 0.5*(600 - 480)*(100 - 80) = $1200.